HELLP!!!!!!!!!! How do you find the height of a similiar triangle???? I have not seen this since 1985!!! One triangle in C A B with the ht being 5 ft and angle C B being 9 ft. The other triangle is F D E with F E being 11 ft.I need to determine the Height of FG from DEF.I would like to try and figure it out myself but I don't know how to and the book for this class (HA-HA) is useless.There are no examples in this book!I have been up all night trying to figure this out!!!
Answer by stanbon(48535)

How do you find the height of a similiar triangle???? I have not seen this since 1985!!! One triangle in C A B with the ht being 5 ft and angle C B being 9 ft. The other triangle is F D E with F E being 11 ft.I need to determine the Height of FG from DEF.I would like to try and figure it out myself but I don't know how to and the book for this class (HA-HA) is useless.There are no examples in this book!I have been up all night trying to figure this out!!!
-----------
I cannot tell from your description what the corresponding sides are.
---
If you said ABC is similar to DEF I would know that AB corresponds to DE
and BC corresponds to EF because of the position in the description.
Then if you said that AB was 5 and DE was 8 and you asked what EF is
if BC is 7, I would set up a proportion and solve for EF as follows:
EF/BC = DE/AB
EF/7 = 8/5
EF = 7(8/5)
EF = 56/5 = 11 1/5 or 11.2
===============================
The heights would also follow the same proportion.
===============================
Cheers,
Stan H.